Show that y= c1ex + c2e2x is the general solution of y’’-3y’+2y=0 on any interval, and find the particular solution for which y(0)=-1 and y’(0)=1


1
Expert's answer
2022-02-21T16:32:27-0500

Given:

"y= c_1e^x + c_2e^{2x}"


"y'= c_1e^x + 2c_2e^{2x}"

"y''= c_1e^x + 4c_2e^{2x}"

Then


"y''-3y'+2y=\\
=(c_1e^x + 4c_2e^{2x})-3( c_1e^x + 2c_2e^{2x})+\\
+2( c_1e^x + c_2e^{2x})=0"

The initial conditions gives

"y(0)= c_1 + c_2=-1"

"y'(0)= c_1 +2 c_2=0"

Finally

"y=-2e^x+e^{2x}"